# Multitask and Multimodal Neural Tuning for Large Models

### v1

@@ Abstract
In recent years, large-scale multimodal models have demonstrated impressive capabilities across various domains. However, enabling these models to effectively perform multiple multimodal tasks simultaneously remains a significant challenge. To address this, we introduce a novel tuning method called neural tuning, designed to handle diverse multimodal tasks concurrently, including reasoning segmentation, referring segmentation, image captioning, and text-to-image generation. Neural tuning emulates sparse distributed representation in human brain, where only specific subsets of neurons are activated for each task. Additionally, we present a new benchmark, MMUD, where each sample is annotated with multiple task labels. By applying neural tuning to pretrained large models on the MMUD benchmark, we achieve simultaneous task handling in a streamlined and efficient manner. All models, code, and datasets will be publicly available after publication, facilitating further research and development in this field.
Multitask and Multimodal Neural Tuning for Large Models
Hao Sun1, Yu Song2, Jihong Hu2, Yen-Wei Chen2, Lanfen Lin1, 1College of Computer Science and Technology, Zhejiang University 2College of Information Science and Engineering, Ritsumeikan University Email1: {sunhaoxx, llf}@zju.edu.cn Email2: {yusong@fc, gr0609ik@ed, chen@is}.ritsumei.ac.jp
@@ 1 Introduction
With the significant progress of large language models (LLMs), researchers have proposed combining large models from multiple modalities to handle multimodal tasks, such as referring segmentation and text-to-image synthesis Ren et al. (2023); Ho et al. (2020). However, these methods primarily focus on individual tasks by designing sophisticated structures or strategies. While this approach can enhance performance in specific areas, the specialized architectural design may limit their applicability to other tasks and domains. Although it is possible to expand the capabilities of LLMs to handle more tasks by developing more intricate structures or tuning strategies, this approach faces scalability challenges.
Figure 1: The overview of our proposed multitask neural tuning. All inputs and outputs are token-based, encompassing both texts and images. The model generates specific tokens for different tasks, such as <OBJ> for segmentation and <GEN> for text-to-image generation. During tuning, a new sparse task network is introduced to emulate SDR and provide task guidance for pretrained LLMs. The entire LLM remains frozen, with only the newly introduced parameters being tunable.
Humans inherently possess multitask learning abilities and can easily adapt to new tasks. One reason for this is the principle known as Sparse Distributed Representation (SDR) or the Sparse Coding Hypothesis (SCH), which describes a way of representing data where only a small fraction of neurons are active at any given time Ahmad and Hawkins (2016); Spanne and Jörntell (2015). This principle is thought to contribute to efficient information processing and the capacity of unique representations in the brain. Inspired by these findings in brain science, we propose a new tuning strategy for multimodal multitask learning, termed neural tuning (as illustrated in Figure 1). Our neural tuning strategy introduces two novel concepts. Firstly, we formulate different tasks in an all-in-token manner, allowing new tasks to be introduced by adding new task tokens, thereby ensuring scalability for additional tasks or modalities. Secondly, we introduce a new sparse task network to emulate SDR for multitask tuning, where only a subset of neurons is activated for each task. The neurons activated by different tasks include both intersecting and unique parts: the intersecting parts represent common information required for various tasks, while the unique parts represent the distinct characteristics of each task. In the design of neural tuning, we avoid additional cross-attention mechanisms and complex decoders to ensure simplicity and efficiency.
However, there are few datasets designed for multimodal multitask learning scenarios, particularly those involving reasoning segmentation tasks where multiple objects can only be segmented through complex reasoning based on both images and text. To address this gap and advance research in this area, we propose a new dataset called MMUD. Constructed using GPT-4 and further annotated by humans, MMUD contains over 36,000 samples. Each sample includes an image with an annotated content caption, a complex reasoning question-answer pair, and referring segmentation masks. We fine-tune pretrained LLMs with neural tuning on MMUD for four tasks: vanilla referring segmentation, reasoning segmentation, image captioning, and text-to-image generation. Figure 1 the demonstrates the pipeline of our proposed multitask neural tuning and the results show that we can achieve state-of-the-art performance.
In summary, our contributions are threefold:
@@
  • We propose neural tuning, a novel framework that unifies diverse multimodal tasks through a concise all-in-token methodology. This innovative approach simplifies the integration of new tasks by merely introducing new task-specific tokens, thus enhancing the flexibility and scalability of large multimodal models.
  • Our method incorporates a sparse task tuning network designed to emulate SDR for various tasks. This network adaptively activates specific neurons to efficiently manage and differentiate between multiple tasks, improving both the precision and adaptability of task performance.
  • We introduce a comprehensive new multimodal benchmark, MMUD, wherein each sample is meticulously annotated for multiple tasks. By fine-tuning models on MMUD, we demonstrate the capability to achieve superior multitask and multimodal task processing performance.
@@ 2 Related Works
In this section, we analysis some recent works related to our works, including multimodal tuning, referring segmentation, and text-to-image synthesis.
@@
2.1 Multimodal Tuning for Large Models
Recently, many multimodal tuning strategies have been proposed based on language models, aiming to equip LLMs with multimodal perception capabilities. For instance, Flamingo Alayrac et al. (2022) employs a frozen image encoder and integrates multimodal signals using gated cross-attention, demonstrating the potential of LLMs to understand multiple modalities. Then BLIP-2 Li et al. (2023a) introduces the Q-Former module, a sophisticated structure designed to align image and text embeddings by extracting common semantics between modalities. This approach has gained widespread adoption due to its effectiveness. Additionally, FROMAGe Koh et al. (2023) employs linear transformations to ground text features in the visual domain, facilitating seamless translation between text and images. Nevertheless, recent vision-language LLMs (LVLMs) have shown great capability in multimodal understanding, such as LLaVA Liu et al. (2023) and Qwen-VL Bai et al. (2023). Our proposed neural tuning strategy is also based on multimodal tuning of LLMs; however, it places greater emphasis on multitask learning in a more elegant and concise manner (especially for reasoning segmentation under complex scenarios) .
2.2 Referring Segmentation and Reasoning Segmentation
Referring segmentation aims to segment objects in an image based on instructive texts. In earlier years, researchers typically employed a text encoder along with a U-Net-like Ronneberger et al. (2015) vision backbone to generate segmentation masks, such as LAVT Yang et al. (2022) and SLViT Ouyang et al. (2023). With the rapid development of large models, simple referring segmentation tasks cannot adequately verify the multimodal reasoning capabilities of these models. To address this challenge, complex reasoning segmentation is proposed which involve asking relatively complex questions about images that require reasoning based on the image content to respond. When the model answers the question, it should also segment the objects involved. For example, Lai et al. (2023) first designed a new LISA model and introduced a new token for segmentation. PixelLM Ren et al. (2023) was then proposed with a newly designed codebook for multi-instance segmentation.
2.3 Text-to-Image Synthesis
There are two text-to-image synthesis approaches highly related to our work: vector quantized generative adversarial network (VQGAN) related methods Esser et al. (2021) and diffusion-based methods Ho et al. (2020); Song et al. (2020). VQGAN aims to map images into a discrete latent space while diffusion models simulates a diffusion process, where data is gradually transformed from a simple prior distribution (like Gaussian noise) to the complex target distribution. State-of-the-art text-to-image models such as GLIDE Nichol and Dhariwal (2021), DALL-E Ramesh et al. (2021), and Imagen Saharia et al. (2022) leverage advanced techniques in VQGAN and diffusion models to generate high-quality, diverse images. In our work, we primarily employ VQGAN for synthesizing images, but we also explore the potential of combining our approach with diffusion networks for image generation.
@@ 3 Method
The overall pipeline of our proposed neural tuning is illustrated in Figure 1. We introduce a new sparse task network into pretrained models, serving as the SDR or SCH. The entire model takes image and text inputs and produces corresponding outputs for different tasks, such as multi-instance referring segmentation (<OBJ>) and image generation (<GEN>). All input and output data are in token format. During the process, the sparse task network activates only a certain percentage of neurons for different tasks. In this section, we first introduce the detailed design of our proposed neural tuning and then provide some details for multitask training.
@@ 3.1 Neural Tuning for Large Models
For a multimodal input comprising an image and a sentence, we first embed the text into $I_{txt}=\{I_{txt}^{l}\}_{l=1}^{L_{t}}$, where $L_{t}$ represents the textual length. Subsequently, we utilize a pretrained vision encoder to extract features from the image, resulting in $I_{img}=\{I_{img}^{l}\}_{l=1}^{L_{i}}$, with $L_{i}$ patches. To integrate the multimodal input, we concatenate the text embeddings and image features (if the task requires visual input) to form inputs for the pretrained LLM: $I=[I_{img};I_{txt}]$.
To fine-tune the pretrained LLMs, we introduce a sparse task network. It is parallel to pretrained LLMs but gets linked in each layer for task guidance. For each tuning layer, it maintains a learnable vector named the Sparse Vector and the details are shown in Figure 2(a). In the $k$-th layer of the LLM before self-attention, we first project the hidden embeddings $H$ into a subspace to obtain the sparse vector:
@@
where $Z_{k-1}^{u}$ is the sparse vector from the last layer. This approach allows the sparse vector to not only receive information from the previous layer but also be aware of the current layer’s information. To enable sparse distributed representation during model flow, only a subset of the sparse vector’s nodes are activated for various tasks. We first randomly sample an activation rate $r$ from a normal distribution $p(r)=N(r;\beta,(0.1\beta)^{2})$, where $\beta\in(0,1)$ is a predefined hyperparameter [^1] More activation patterns are discussed in Appendix C. Then, we activate corresponding neurons as follows:
@@
In the LLM output stage, we introduced new tokens for multimodal tasks alongside the original textual tokens. For example, we introduced <OBJ> tokens for segmentation and <GLB> tokens for text-to-image synthesis. Different task-specific tokens are fed into their respective decoders. This approach allows us to unify the input and output of all tasks into an all-in-token manner, making it convenient to add more tasks and modalities.
@@ 3.2 Multitask Training
As different tasks are unified into an all-in-one token manner, neural tuning can be conducted using a simple causal language modeling approach. Cross-entropy loss is employed as the loss function ($\mathcal{L}_{txt}$) for next-token prediction.
For segmentation-related tasks, the embeddings corresponding to segmentation tokens <OBJ>, $H_{seg}$, are collected and fed to a lightweight decoder for corresponding mask generation:
@@
where $W_{seg}$ represents the learnable parameters in the mask decoder. Different from previous methods Lai et al. (2023), which employ pretrained SAM Kirillov et al. (2023) for predicting masks, the designed decoder is much lighter. Since the models can generate multiple segmentation tokens at a time, we can easily perform multi-instance segmentation. Following previous works Lai et al. (2023); Ren et al. (2023), DICE loss Milletari et al. (2016) is employed in our framework to guide the segmentation tasks:
@@
In text-to-image synthesis tasks, we employ pretrained VQGAN generators Esser et al. (2021); Wang et al. (2022) to synthesize images. To generate a sequence of indices for VQGAN to produce images, we train a conditional transformer to predict the indices in an autoregressive manner:
@@
where $\mathbf{s}$ is the sequence of indices for VQGAN to generate images and $c$ is the condition that controls the contents of the images. To simplify the tuning process, we first pretrain the conditional transformer on MS-COCO Lin et al. (2014) and employ the image embeddings from the pretrained CLIP model Radford et al. (2021) as the condition. After training, we can generate images via the CLIP embeddings. Therefore, in neural tuning, we simply align the embeddings of the synthesis tokens ($H_{gen}$) to the CLIP embeddings ($H_{clip}$). The mean squared error is employed to perform the alignment:

### v2

# One Framework to Rule Them All: Unifying Multimodal Tasks with LLM Neural-Tuning

@@ Abstract
Large-scale models have exhibited remarkable capabilities across diverse domains, including automated medical services and intelligent customer support. However, as most large models are trained on single-modality corpora, enabling them to effectively process and understand multimodal signals remains a significant challenge. Current research often focuses on designing task-specific or scenario-specific tuning strategies, which limits the scalability and versatility. To address this limitation, we propose a unified framework that concurrently handles multiple tasks and modalities. In this framework, all modalities and tasks are represented as unified tokens and trained using a single, consistent approach. To enable efficient multitask processing, we introduce a novel tuning strategy termed neural tuning, inspired by the concept of sparse distributed representation in the human brain, where only specific subsets of neurons are activated for each task. Furthermore, to advance research in multimodal and multitask learning, we present a new benchmark, MMUD, which includes samples annotated with multiple task labels spanning reasoning segmentation, referring segmentation, image captioning, and text-to-image generation. By applying neural tuning to pretrained large models on the MMUD benchmark, we demonstrate the ability to handle multiple tasks simultaneously in a streamlined and efficient manner. All models, code, and datasets will be released publicly upon publication, fostering further research and innovation in this field.
keywords:
Multimodal learning, Large language models, Pretrained model tuning, Referring segmentation, Complex Segmentation, Image generation
journal: ArXiv
[…]
[label1]organization=College of Computer Science and Technology, Zhejiang University, city=Hangzhou, state=Zhejiang, country=China
[…]
[label2]organization=College of Information Science and Engineering, Ritsumeikan University, city=Ibaraki, state=Osaka, country=Japan
@@ 1 Introduction
Recently, the rapid advancements in deep learning and hardware computing power have propelled the development of large-scale models, achieving significant breakthroughs in applications such as intelligent customer support and autonomous driving. The remarkable success of these models can be attributed to their large-scale architectures and extensive training data, which substantially enhance their contextual understanding and complex reasoning capabilities. However, most large-scale models are pretrained on single-modality corpora due to the challenges associated with acquiring large-scale multimodal datasets and the computational limitations of current hardware. To address this, recent research has explored enabling large language models (LLMs) to process multimodal data through fine-tuning. While these methods improve performance on specific tasks, they predominantly rely on task-specific architectures or tuning strategies (such as referring segmentation and image-text classification [1, 2, 3]), which significantly hinders their scalability and versatility in handling multiple tasks concurrently. Although expanding the capabilities of LLMs through more intricate architectures or tuning strategies is possible, this approach imposes considerable complexity and overhead, limiting the feasibility of extending such models to accommodate additional tasks or datasets.
In revisiting the human cognitive process, which serves as the inspiration for artificial intelligence, we observe that humans inherently excel at multitask learning and effortlessly adapt to new tasks. A key factor underpinning this ability is the principle of Sparse Distributed Representation (SDR), or the Sparse Coding Hypothesis (SCH), which posits that information is represented in a way where only a small fraction of neurons are active at any given time [4, 5]. This principle is believed to enhance the efficiency of information processing and enable the brain to form unique and robust representations of complex data. Motivated by these findings in neuroscience, we propose a unified framework for multitask and multimodal learning in large language models (LLMs), as illustrated in Figure 1. Our framework introduces two key innovations:
Unified Tokenization for Multimodal Multitask Learning. We formulate all tasks, as well as their multimodal data inputs and outputs, into a unified token-based representation. Unlike previous approaches that rely on cross-attention mechanisms to facilitate multimodal interactions [6, 7], our framework directly feeds the tokenized inputs into the model, leveraging the pretrained model’s self-attention mechanism [8]. This allows the model to compute relationships across and within modalities in a holistic manner, enhancing its ability to understand inter- and intra-modal dependencies. By unifying tasks and modalities in this manner, we train all multimodal tasks simultaneously, enabling the model to capture shared patterns across different input data. This approach not only simplifies the tuning process but also mitigates task-specific biases, as the model learns to generalize from diverse input-output mappings. Consequently, our method allows pretrained models to handle multiple tasks with a single causal language modeling objective, significantly reducing computational overhead compared to prior methods.
Neural Tuning Strategy. To further improve the efficiency and adaptability of multimodal multitask learning, we introduce neural tuning, a novel tuning strategy inspired by SDR. In this strategy, only a subset of neurons is activated for each task, mimicking the sparse activation patterns observed in the human brain. Neurons activated by different tasks consist of two components: shared neurons, which capture common features across tasks, and task-specific neurons, which encode unique characteristics for individual tasks. This design not only enhances multitask learning efficiency but also reduces computational costs by selectively activating neurons. Furthermore, by avoiding additional cross-attention mechanisms and complex decoders, our framework ensures simplicity and computational efficiency while maintaining strong performance across tasks.
In summary, our framework seamlessly integrates multimodal and multitask learning into a unified architecture inspired by human cognitive principles, enabling efficient task handling with reduced computational requirements. Despite these advancements, there is currently a lack of datasets specifically designed for multimodal multitask learning, particularly for challenging tasks like reasoning segmentation, where multiple objects can only be segmented through complex reasoning that combines both image and textual information. To address this gap and foster progress in this domain, we introduce a novel dataset, MMUD (Multimodal Understanding Dataset). MMUD is constructed using GPT-4 to generate initial annotations, which are subsequently refined and verified through human annotation. The dataset comprises over 36,000 samples, each containing an image paired with a detailed content caption, a complex reasoning question-answer pair, and referring segmentation masks that align with object words in intricate descriptions. MMUD is explicitly designed to support diverse multimodal tasks that demand advanced reasoning and multimodal understanding. A comprehensive description of the dataset and its construction process is provided in Section 4. To demonstrate the capabilities of our framework, we fine-tune pretrained LLMs on MMUD for four distinct tasks: vanilla referring segmentation, reasoning segmentation, image captioning, and text-to-image generation. Figure 1 illustrates the pipeline of our proposed method, incorporating neural tuning. Experimental results indicate that our approach achieves state-of-the-art performance across these tasks, highlighting the effectiveness and generalizability of the framework.
Our contributions are threefold:
  • 1.
@@
    We propose a novel framework that unifies diverse multimodal tasks using a concise all-in-token methodology. This approach simplifies the integration of new tasks by requiring only the introduction of task-specific tokens, significantly enhancing the flexibility and scalability of large multimodal models.
  • 2.
    We introduce neural tuning, a sparse task-tuning strategy inspired by Sparse Distributed Representation (SDR). This approach adaptively activates specific subsets of neurons for different tasks, enabling efficient multitask management while enhancing precision and adaptability across tasks.
  • 3.
    We present a new multimodal benchmark, MMUD, which includes meticulously annotated samples designed for multiple tasks, such as reasoning segmentation, image captioning, and text-to-image generation. By fine-tuning models on MMUD using our proposed framework, we demonstrate superior multitask and multimodal processing capabilities, setting a new state-of-the-art in performance.
The remainder of this paper is organized as follows. In Section 2, we review related works, focusing on advancements in multimodal learning, tuning strategies, and their applications in large language models. Section 3 presents our proposed unified framework, detailing its innovative all-in-token methodology and the neural tuning strategy for efficient multitask and multimodal learning. In Section 4, we introduce the MMUD dataset, describing its construction, annotation process, and utility for evaluating multitask and multimodal learning. Section 5 illustrates the experimental results and provides a comprehensive analysis of the proposed framework’s performance across various tasks. Finally, Section 6 concludes the paper by summarizing our contributions and discussing potential directions for future research.
Figure 1: The overview of our proposed multitask and multimodal tuning framework. All inputs and outputs are token-based, encompassing both texts and images. The model generates specific tokens for different tasks, such as $<$OBJ$>$ for segmentation and $<$GEN$>$ for text-to-image generation. During tuning, a new sparse task network is introduced to emulate SDR and provide task guidance for pretrained LLMs. The entire LLM remains frozen, with only the newly introduced parameters being tunable.
@@ 2 Related Works
In this section, we analysis some recent or foundational works related to our method, including multimodal tuning, referring segmentation, and text-to-image synthesis.
2.1 Multimodal Tuning for Large Models
Recently, large models have demonstrated remarkable potential across diverse domains, not only in practical applications but also as tools for advancing scientific inquiry. However, training such models directly on multimodal data is often infeasible due to significant hardware and data constraints. To address this, various multimodal tuning strategies have emerged, primarily leveraging pre-trained language models to endow them with multimodal perception capabilities. The core concept of these approaches is to preserve the original model parameters by keeping them frozen, while adapting newly introduced components through lightweight methods such as adapters or specialized architectures. For example, Flamingo [9] retains a frozen image encoder and incorporates multimodal information via gated cross-attention, showcasing the ability of large language models (LLMs) to process multiple modalities. BLIP-2 [10], on the other hand, introduces the Q-Former module, a sophisticated mechanism designed to align image and text embeddings by extracting shared semantic representations, which has proven highly effective and widely adopted. Similarly, FROMAGe [11] employs linear transformations to bridge textual and visual features, enabling seamless translation between these domains. More recently, vision-language LLMs (LVLMs) such as LLaVA [12] and Qwen-VL [13] have demonstrated exceptional multimodal reasoning and understanding capabilities. Our proposed framework and the novel neural tuning strategy build on the foundations of multimodal tuning for LLMs. However, our approach emphasizes multitask learning in a more elegant and concise manner, particularly excelling in reasoning and segmentation tasks under complex scenarios.
2.2 Referring Segmentation and Reasoning Segmentation
Among multimodal understanding tasks, referring segmentation is a foundational challenge that involves segmenting objects in an image based on textual instructions. This task evaluates a model’s ability to establish connections between fine-grained visual details and linguistic cues, making it a crucial benchmark for assessing multimodal understanding. Earlier approaches typically utilized a text encoder alongside a U-Net-like [14] vision backbone to produce segmentation masks, as exemplified by methods such as LAVT [15] and SLViT [6]. With the advent of large models, however, referring segmentation has become relatively trivial, as simple textual instructions fail to challenge the advanced reasoning capabilities of these models. To address this limitation, the concept of complex reasoning segmentation was introduced. This paradigm extends beyond basic segmentation by incorporating complex questions about image content that require sophisticated reasoning to answer. The output is not limited to segmentation masks but also includes detailed textual responses to these complex queries. Furthermore, the model must identify and segment the objects relevant to its answers. For example, LISA [16] introduced a novel architecture with specialized tokens for segmentation, while PixelLM [2] proposed a newly designed codebook to support multi-instance segmentation. In our work, reasoning segmentation is a pivotal task for evaluating the multimodal reasoning and processing capabilities of our framework.
2.3 Text-to-Image Synthesis
For text-to-image synthesis, there are generally two approaches highly related to our work: vector quantized generative adversarial network (VQGAN) related methods [17] and diffusion-based methods [3, 18]. VQGAN aims to map images into a discrete latent space while diffusion models simulates a diffusion process, where data is gradually transformed from a simple prior distribution (like Gaussian noise) to the complex target distribution. State-of-the-art text-to-image models such as GLIDE [19], DALL-E [20], and Imagen [21] leverage advanced techniques in VQGAN and diffusion models to generate high-quality, diverse images. In our work, we primarily employ VQGAN for synthesizing images, but we also explore the potential of combining our approach with diffusion networks for image generation.
@@ 3 Method
In this section, we first describe the detailed design of our framework with neural tuning, and then outline the specifics of multitask training.
3.1 All-in-token Multimodal Paradigm
The overall pipeline of our proposed neural tuning framework is illustrated in Figure 1. The inputs from the different modalities are tokenized and processed concurrently, emulating the ability of humans to seamlessly handle multiple tasks. We also introduce a novel sparse task network into pretrained models, which functions as either the Sparse Distributed Representation (SDR) or the Sparse Coding Hypothesis (SCH), through which only a subset of neurons are activated, tailored to the specific task at hand. This paradigm enables the model to take both image and text inputs, generating task-specific outputs such as multi-instance referring segmentation ($<$OBJ$>$) and image generation ($<$GEN$>$).
Specifically, for a multimodal input consisting of images and sentences, we first embed the text into $I_{txt}=\{I_{txt}^{l}\}_{l=1}^{L_{t}}$, where $L_{t}$ denotes the length of the text. Next, we use a pretrained vision encoder to extract features from the image, resulting in $I_{img}=\{I_{img}^{l}\}_{l=1}^{L_{i}}$ where $L_{i}$ representing the number of image patches. To integrate the multimodal input, we concatenate the text embeddings and image features (when the task requires visual input) to form the final input for the pretrained large language model (LLM): $I=[I_{img};I_{txt}]$. In this scheme, the textual and visual modalities interact through the LLM’s self-attention mechanism, where the output $O$ is computed as $O=SoftMax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ with the Query($Q$), Key($K$), and Value($V$) derived from the concatenated multimodal input $I$. Unlike prior works that employ cross-attention for modality interaction [6, 15], where $Q$, $K$, and $V$are derived from different modalities, our self-attention approach offers several advantages. Not only does it simplify the architecture by avoiding explicit cross-modal mappings, but it also enables the model to compute both inter- and intra-modal relationships. This dual interaction within the self-attention framework enhances the model’s ability to understand and fuse multimodal information more effectively.
In the output stage of the LLM, we introduce new task-specific tokens alongside the original textual tokens to handle multimodal tasks. For example, we incorporate $<$OBJ$>$ tokens for segmentation tasks and $<$GLB$>$ tokens for text-to-image synthesis. These task-specific tokens are then fed into their respective decoders, enabling the model to generate the appropriate outputs for each task. This approach unifies the input and output format across all tasks into an all-in-token scheme, simplifying the integration of additional tasks and modalities. As a result, the system becomes more flexible and scalable, allowing for easy extension to new multimodal tasks without significant architectural changes.
@@ 3.2 Neural Tuning for Large Models
To fine-tune the pretrained LLMs, we introduce a sparse task network behaving like SDR. It is parallel to pretrained LLMs but gets linked in each layer for task guidance. For each tuning layer, it maintains a learnable vector named the Sparse Vector and the details are shown in Figure 2(a). In the $k$-th layer of the LLM before self-attention, we first project the hidden embeddings $H$ into a subspace to obtain the sparse vector:
@@
where $Z_{k-1}^{u}$ is the sparse vector from the last layer. This approach allows the sparse vector to not only receive information from the previous layer but also be aware of the current layer’s information. To enable sparse distributed representation during model flow, only a subset of the sparse vector’s nodes are activated for various tasks. We first randomly sample an activation rate $r$ from a normal distribution $p(r)=N(r;\beta,(0.1\beta)^{2})$, where $\beta\in(0,1)$ is a predefined hyperparameter [^1] More activation patterns are discussed in Section 5.5. Then, we activate corresponding neurons as follows:
@@ 3.3 Multitask Training
As different tasks are unified into an all-in-one token manner, the tuning procedure can be conducted using a simple causal language modeling approach. Cross-entropy loss is employed as the loss function ($\mathcal{L}_{txt}$) for next-token prediction:
$\mathcal{L}_{txt}=-\sum_{t}\log\hat{P}(x_{t}|x_{i};i<t),$ (6)
where $\hat{P}(x_{t}|x_{i};i<t)$ is the predicted probability for the token $x_{t}$ based on the context of all previous tokens.
For segmentation-related tasks, the embeddings corresponding to object tokens $<$OBJ$>$, $H_{seg}$, are collected and fed to a lightweight decoder for corresponding mask generation:
@@
where $W_{seg}$ represents the learnable parameters in the mask decoder. Different from previous methods [16], which employ pretrained SAM [22] for predicting masks, our designed decoder is much lighter. The mask decoder consists of three convolutional layers for upscaling, each followed by three self-attention layers for mask decoding. Since the models can generate multiple segmentation tokens at a time, we can easily perform multi-instance segmentation. Following previous works [16, 2], DICE loss [23] is employed in our framework to guide the segmentation tasks:
@@
In text-to-image synthesis tasks, we employ pretrained VQGAN generators [17, 24] to synthesize images. To generate a sequence of indices for VQGAN to produce images, we train a conditional transformer to predict the indices in an autoregressive manner:
@@
where $\mathbf{s}$ is the sequence of indices for VQGAN to generate images and $c$ is the condition that controls the contents of the images. To simplify the tuning process, we first pretrain the conditional transformer on MS-COCO [25] and employ the image embeddings from the pretrained CLIP model [26] as the condition. After training, we can generate images via the CLIP embeddings. Therefore, in neural tuning, we simply align the embeddings of the synthesis tokens ($H_{gen}$) to the CLIP embeddings ($H_{clip}$). The mean squared error is employed to perform the alignment: